IDENTIFICATION OF THE LOWEST-LYING BENDING VIBRATIONAL LEVELS OF THE Ã ^{1}A_{u} STATE OF ACETYLENE, C_{2}H_{2}

\underline{ANTHONY J. MERER}, Institute of Atomic and Molecular Sciences, Academia Sinica, P.O. Box 23-166, Taipei, Taiwan 10617, Department of Chemistry, University of British Columbia, 6174 University Boulevard, Vancouver, B.C., Canada V6T 1Z3; NAMI YAMAKITA, Department of Chemical and Biological Sciences, Japan Women’s University, Mejirodai, Bunkyo-ku, Tokyo 112-8681, Japan; ADAM H. STEEVES, HANS A. BECHTEL and ROBERT. W. FIELD, Department of Chemistry, Massachusetts Institute of Technology, Cambridge, MA 02139.

Acetylene is linear in its ground state, but becomes trans-bent in its Ã ^{1}A_{u} state. The vibrational structure of the Ã ← X̃ system consists^{a} of long progressions in the ν′_{3} (trans-bending) vibration based on the origin and the ν′_{2} (C-C stretch) fundamental. These progressions become irregular at a vibrational energy of about 5000 cm^{−1}, with many extra bands appearing^{b}. Spectra of jet-cooled acetylene showed recently^{c} that extra bands can also be found at much lower energy. Two of the extra levels were identified as the a_{g} combinations 1^{1}3^{1} and 2^{2} 3^{1}, but the other levels are combinations involving the two low frequency bending vibrations, ν_{4} (torsion) and ν_{6} (in-plane bending). Two types of high resolution spectra have been recorded in order to assign these extra levels. Infrared-ultraviolet double resonance spectra via the X̃, ν_{3} level have identified large numbers of ungerade vibrational levels at energies of 2000-7000 cm^{−1}, while high sensitivity laser-induced fluorescence spectra of jet-cooled acetylene have given the positions of many low-lying gerade levels. The low-lying bending levels are affected by strong a-axis Coriolis coupling, which distorts their K -structures considerably. In addition, there is very strong Darling-Dennison resonance between ν_{4} and ν_{6}, with k_{4466} ≈ 50 cm^{−1}. Nevertheless a good fit to the structures of the pure bending polyads can be obtained, allowing for both the Darling-Dennison resonance and the Coriolis coupling.

aJ.K.G. Watson, M. Herman, J.C. Van Craen and R. Colin, J. Mol. Spectrosc., 95, 101 (1982).

bJ.C. Van Craen, M. Herman, R. Colin and J.K.G. Watson, J. Mol. Spectrosc., 111, 185 (1985).

cA.J. Merer, N. Yamakita, S.Tsuchiya, J.F. Stanton, Z. Duan and R.W. Field, Mol. Phys., 101, 663 (2003).